OpenStudy (anonymous):

(1.83*10^6) /( 1.2 * 10^4)Express using scientific notation retain 3 decimals

OpenStudy (anonymous):

Plug it into your calculator, you should get \[1.525\times10^2\]

OpenStudy (anonymous):

This is retaining 3 decimals

OpenStudy (unklerhaukus):

\[(1.83\times10^6) \div( 1.2 \times 10^4)\] \[=\frac{1.83}{1.2}\frac {10^6}{10^4}=\frac {183}{1.2}\times10^{6-4}\]

OpenStudy (anonymous):

Yes, that's fine as well, if that's what they want as an answer

OpenStudy (unklerhaukus):

i was just ring to show the method to use if you dont have a calculator

OpenStudy (unklerhaukus):

trying*

OpenStudy (anonymous):

:) cool